OpenStudy (anonymous):

trigonometric equation sin^2x+cos2x-cosx=0

myininaya (myininaya):

write your equation just in terms of cos(x)

myininaya (myininaya):

tell me though what happens when you apply the double angle identity for cosine? \[\cos(2x)=\cos^2(x)-\sin^2(x) \text{ by double angle identity}\]

OpenStudy (anonymous):

I got sin^2x+(cos^2x-sin^2x)-cosx=0